In the following problems, solve the given initial value problem using the method of Laplace transforms (a) y – 7y + 10y =
Answer #1

. On both sides of equation. (a) Apply haplace transton L{y } - #2 {y } + 102{y} - 9w{ costy + 728 Sht} Sa Lly)- syco) 10)-7Put and s=0 A=8, B=-4 639) 0++ (-2)(-5) (1) $+ + - 4 -5 D f 45 0 De-32 538 + Put A-6, B=-4,D=8 and S= 1 (5-39 )(19+1)+9+7 -4